# Reconfigurable routing in data center networks

### v1

@@ Abstract
The Reconfigurable Routing Problem (RRP) in hybrid networks is, in short, the problem of finding settings for optical switches augmenting a static network so as to achieve optimal delivery of some given workload. The problem has previously been studied in various scenarios with both tractable and NP-hardness results obtained. However, the data center and interconnection networks to which the problem is most relevant are almost always such that the static network is highly structured whereas all previous results assume that the static network can be arbitrary (which makes existing computational hardness results less technologically relevant and also easier to obtain). In this paper, and for the first time, we prove various intractability results for RRP where the underlying static network is highly structured, for example consisting of a hypercube, and also extend some existing tractability results.
@@ 1 Introduction
The rapid growth of cloud computing applications has induced demand for new technologies to optimize the performance of data center networks dealing with ever-larger workloads. The data center topology design problem (that of finding efficient data center topologies) has been studied extensively and resulted in myriad designs (see, e.g., CGC16 ). Advances in hardware, such as optical switches reconfigurable in milli- to micro-seconds, have enabled the development of reconfigurable topologies (see, e.g., HFS21 ). These topologies can adjust in response to demand (demand-aware reconfigurable topologies) or vary configurations over time according to a fixed protocol (demand-oblivious reconfigurable topologies; see, e.g., AS19 ). So-called hybrid data center networks are a combination of a static topology consisting of, for example, electrical switches, and a demand-aware reconfigurable topology implemented, for example, with optical circuit switches or free space optics (see, e.g., FS19 ).
@@
The hybrid network paradigm combines the robustness guarantees of static networks with the ability of demand-aware reconfigurable networks to serve large workloads at very low cost. Of particular interest to us is the question of how the reconfigurable (optical) portion of the network should be configured for some demand pattern, formalized by Foerster et al. FGS18 as the Reconfigurable Routing Problem (RRP): in short, we are given a static network, augmented with some optical switches, and a workload and we wish to find settings for the switches which result in an optimal delivery of the workload (full definitions follow). Prior work FFSV20 ; FGS18 ; FPS19 has shown that RRP is computationally hard but only when the static network is allowed to be arbitrary.
@@
Data center networks, and interconnection networks in general, typically have highly structured topologies. For example: the popular switch-centric data centre network Fat-Tree ALV08 is derived from a folded Clos network; the server-centric data centre network DCell GWT08 is recursively-structured whereby at each level, a graph-theoretic matching of servers is imposed; and the server-centric data centre network BCube GLL09 is recursively-structured with a construction based around a generalized hypercube. Many (but not all) NP-complete problems become tractable when the input is restricted to the graphs providing the communications fabric for data center networks and other interconnection networks; for example, Hamiltonian cycles are often trivial to find in many interconnection networks.
@@
As stated above, previous research on RRP has not investigated the computational complexity of the problem when the static portion of the network is highly structured (as is almost always the case in existing data center topologies). The question of the complexity of RRP for specific network topologies was specifically identified as an area for future work in FFSV20 . In light of this, we establish for the first time hardness results for RRP that apply to various specific families of highly structured static networks such as, for example, the hypercubes. Our constructions are (perhaps not surprisingly) of a much more involved nature than has hitherto been the case.
In Section 2.1, we detail our hybrid network model and define our fundamental reconfiguration routing problem RRP before highlighting some parameterized restrictions on our model and their technological relevance in Section 2.2. In Section 3, we catalogue existing results regarding RRP and explain any variations between the models assumed for these results and our model from Section 2.2. We present our results in Section 4, together with a commentary, and include the proofs of these results in Section 5.
2 Models of reconfigurable networks
We begin by making precise our model of reconfigurable networks and defining the fundamental reconfiguration routing problem before looking at specific restrictions, motivated by technological considerations, that might be imposed on this problem.
2.1 The fundamental reconfiguration problem
@@
Definition 1 (hardware).
@@
A static network $G$ can be abstracted as an undirected graph $G=(V,E)$ so that each static link $(u,v)\in E$ has weight $w(u,v)\in\mathbb{Q}_{+}$ (reflecting a transmission cost) and is incident with internal ports of two distinct nodes of $V$. The number of internal ports of some node $v\in V$ is $\rho_{i}(v)\in\mathbb{N}$ (so, $\rho_{i}(v)$ is the degree of $v$ in the abstracted graph $G$). There is a set $S$ of switches and there are switch links joining switch ports of some switch to external ports of some of the nodes of $V$. Every switch link has weight 00 (we say more about switch link weights momentarily). Every switch $s\in S$ has $\rho_{s}(s)\in\mathbb{N}\setminus\{0,1\}$ switch ports and every node $v\in V$ has $\rho_{e}(v)\in\mathbb{N}$ external ports. In general, the number of external ports of the nodes of $V$ is variable, as is the number of switch ports of the switches of $S$, and it may be the case that there is more than one switch link between a specific node and a specific switch (henceforth, it is assumed that every external port of every node is incident with some switch port and that every switch port of every switch is incident with some external port of a node, by simply ignoring all unused ports). The static network $G$ together with the switches of $S$ is known as the hybrid network $G(S)$. (A hybrid network can be visualized as in Fig. 1.)
Figure 1: A hybrid network.
If we wish to pinpoint some external port of some node (resp. switch port of some switch) then we write $u^{i}$ (resp. $s^{i}$) to denote external port $i$ of node $u$ (resp. switch port $i$ of switch $s$).
Definition 2 (configuring the hardware).
@@
Given a hybrid network $G(S)$ and a switch $s\in S$, a switch matching $N_{s}$ of $s$ is a set of pairs of switch ports of $s$ so that all switch ports involved are distinct. Each switch matching represents an internal setting of the switch and naturally yields a set of pairs of external ports of nodes where all such ports are distinct; we refer to a set of pairs of external ports obtained in this way as a node matching (note that this differs from the standard graph-theoretic notion of a matching). A configuration $N$ is a set of switch matchings $\{N_{s}:s\in S\}$. If the pair $(u^{i},v^{j})$ of external ports is in some node matching resulting from some switch matching $N_{s}$ of some configuration $N$ then we say that there is an $s$-dynamic link joining nodes $u$ and $v$, or a dynamic link for short. We allocate a fixed weight $\mu\in\mathbb{Q}_{+}$ to each internal port-to-port connection in a switch. So, although an $s$-dynamic link is an atomic entity, it can be visualized as consisting of a switch link followed by an internal port-to-port connection in $s$ followed by another switch link. We denote by $G(N)$ the static network $G$ augmented with the dynamic links (each of weight $\mu$) resulting from the configuration $N$ and we call $G(N)$ an augmented network (note that it is possible that $G(N)$ is a multigraph). We sometimes refer to the dynamic links arising from some configuration $N$ as $N$ also. (An augmented network can be visualized on the left in Fig. 2 with the dynamic links abstracted on the right.)
Figure 2: An augmented network $G(N)$ and the abstracted dynamic links.
@@
Definitions 1 and 2 are driven by reconfigurable hardware technology such as optical switches, wireless (beamforming) and free-space optics, all of which establish port-to-port connections, i.e., switch matchings. The survey paper FS19 provides some detail as regards the relationship between the emergent theoretical models and current (optical) technology.
Definition 3 (dealing with workloads).
@@
For some hybrid network $G(S)$, a workload (matrix) $D$ with entries $\{D[u,v]\in\mathbb{Q}_{+}:u,v\in V\}$ provides the intended pairwise node-to-node workloads (each $D[u,u]$ is necessarily 00). Given a configuration $N$ and $u,v\in V$ for which $D[u,v]>0$, we route the corresponding workload via a path in $G(N)$ from $u$ to $v$ in $G(N)$ so that this chosen flow-path $\varphi(u,v)$ has workload cost $D[u,v]\times wt_{G(N)}(\varphi(u,v))$, where the weight $wt_{G(N)}(\varphi(u,v))$ is the sum of the weights of the links of the flow-path $\varphi(u,v)$ (if $G(N)$ has both a static link $(x,y)$ and a dynamic link $(x,y)$ then we need to say which we are using in $\varphi(u,v)$). The total workload cost (of $D$) is defined as
@@
Our aim will be to find a configuration $N$ in some hybrid network $G(S)$ and flow-paths in $G(N)$ for which the total workload cost of some workload matrix $D$ is minimized. In an unrestricted scenario, we would choose any flow-path $\varphi(u,v)$ to be a flow-path of minimum weight from $u$ to $v$ in $G(N)$, the weight of which we denote by $wt_{G(N)}(u,v)$. We also have the analogous concepts $wt_{G}(\varphi(u,v))$ and $wt_{G}(u,v)$ where we work entirely in the static network $G$. Note that we often describe $D$ by a weighted digraph, which we usually call $D^{\prime}$, so that the node set is $V$ and there is an edge $(u,v)$ of weight $w>0$ if, and only if, $D[u,v]=w$. We also refer to some $D[u,v]>0$ as a demand (from $u$ to $v$).
We are now in a position to define our fundamental reconfiguration routing problem RRP.
Definition 4 (problem RRP).
Given a hybrid network $G(S)$, a fixed dynamic link weight $\mu\in\mathbb{Q}_{+}$, a workload matrix $D$ and a value $\kappa\geq 0$, the reconfigurable routing problem RRP is to decide whether there is a configuration $N$ and appropriate work-flow paths in $G(N)$ so that the total workload cost for workload matrix $D$ when every dynamic link has weight $\mu$ is at most $\kappa$.
2.2 Some restrictions and their motivation
There are various natural restrictions we can make on our hybrid networks and on our fundamental problem RRP, motivated by technological considerations such as limitations on the number of available external ports at a node, restrictions on and costs of mixing static (e.g., electrical) and dynamic (e.g., optical) links on a flow-path, and imposed routing policies in relation to workload segregation. These restrictions originated in FFS19 ; FFSV20 ; FGS18 ; FPS19 and we refer he reader to those works for more motivation as regards practical aspects of routing in reconfigurable networks.
Definition 5 (hardware).
Let $G(S)$ be a hybrid network. For some $k\in\mathbb{N}$, we define RRP($\Delta_{S}=k$) to be the problem RRP restricted to instances $(G(S),\mu,D,\kappa)$ where every node $v\in V$ is such that $\rho_{e}(v)\leq k$.
Definition 6 (flow-path policies).
Let $G(S)$ be a hybrid network. We can impose a segregation parameter $\sigma\in\mathbb{N}\cup\{\infty\}$ on the composition of flow-paths so that any flow-path must, by definition, have at most $\sigma$ alterations between static and dynamic links: if $\sigma=0$ then the flow-paths are segregated; and if $\sigma=\infty$ then the flow-paths are non-segregated. For $k\in\mathbb{N}\cup\{\infty\}$, we define RRP($\sigma=k$) to be the problem RRP restricted so that all flow-paths must necessarily have at most $k$ alternations of static and dynamic links.
Segregation, whereby large flows (i.e., elephant flows) are routed along dynamic links and small flows (i.e., mice flows) along static links, is a commonly found routing policy in networks with reconfigurable components. The two extremes are segregated flows, when $\sigma=0$, and non-segregated flows, when $\sigma=\infty$.
Definition 7 (link usage).
Let $G(S)$ be a hybrid network. We can impose a dynamic bound $\delta\in\mathbb{N}\cup\{\infty\}$ (resp. length bound $\lambda\in\mathbb{N}\cup\{\infty\}$) on the composition of flow-paths so that any flow-path must use at most $\delta$ dynamic links (resp. $\lambda$ links) in total. For $k\in\mathbb{N}\cup\{\infty\}$, we define RRP($\delta=k$) (resp. RRP($\lambda=k$)) to be the problem RRP restricted so that all flow-paths necessarily use at most $k$ dynamic links (resp. $k$ links) in total.
We can compose the restrictions in Definitions 5, 6 and 7 to obtain, for example, the problem RRP($\Delta_{S}=k_{1},\sigma=k_{2},\delta=k_{3},\lambda=k_{4}$) whose definition is obvious (given specific values for each $k_{i}$). Furthermore, we use RRP($\sigma\geq k$) for some fixed $k$ to describe results which hold for RRP($\sigma=k^{\prime}$), for any $k^{\prime}\geq k$. Note that a proof that RRP($\sigma\geq k^{\prime}$) is NP-hard is a stronger statement than saying RRP($\sigma=k^{\prime}$) is NP-hard as the former implies the latter but not necessarily vice versa.
3 Existing results
We begin with an overview of existing results before providing consequent motivation for our upcoming analysis.
3.1 Hardness and tractability results
The following results were established in FFS19 ; FFSV20 ; FGS18 ; FPS19 and the reader is referred to these papers for more details. The paper FS19 provides a wider survey of the relationship between technological aspects of reconfigurable data centre networks and the theoretical framework.
It was proven in FGS18 ; FPS19 that the problem RRP is NP-complete, even under various restrictions, with additional complexity results in FFS19 ; FFSV20 . However, as we explain now, the models in FFS19 ; FFSV20 ; FGS18 ; FPS19 are different from ours (as presented in Section 2), although some of the results still hold in our model.
  • 1.
    In FGS18 , variable dynamic link weights are allowed; so, whereas for us every dynamic link has a fixed weight $\mu$, in FGS18 the weight of any dynamic link can potentially be any prescribed value. This makes it easier to secure hardness results than in our model (though any feasibility results are stronger than in our model). It is also assumed in FGS18 that every node has at most one external port (though multiple switches are allowed).
  • 2.
    In FPS19 , variable dynamic link weights are again allowed but it is assumed that there is always only one switch with every node having the same number of external ports (possibly more than one).
  • 3.
    In FFS19 ; FFSV20 , variable dynamic link weights are again allowed and there is a consideration of (static and dynamic) uni-directional links.
The hardness results in FGS18 ; FPS19 are as follows. We present these results using the nomenclature of Section 2 and the underlying model is that of Section 2 unless otherwise stated.
Theorem 1 ((FGS18, , Lemma 1)).
The problem RRP($\Delta_{S}=1$, $\sigma\geq 2$, $\delta\geq 1$, $\lambda\geq 3$) is NP-complete. Moreover, we may assume that: all switches have $3$ ports; all link weights of the static network are $1$; the workload matrix is sparse and its entries are all either 00 or $1$; and the weight of any dynamic link is $\mu=1$.
The construction in the proof of Theorem 1 in FGS18 requires multiple $3$-port switches and allows for the case that there may be nodes with no external ports. However, an analogous result is also given for the case when there is only one switch.
A version of Theorem 1 was also proven where $\delta=2$.
Theorem 2 ((FGS18, , Theorem 2)).
The problem RRP($\Delta_{S}=1$, $\sigma\geq 2$, $\delta\geq 2$, $\lambda\geq 3$) is NP-complete. Moreover, we may assume that instances are such that: all link weights of the static network are $1$; there is one switch and every node of the static network is adjacent to the switch; the workload matrix is sparse and entries are all either 00 or $1$; and the weight of any dynamic link is $\mu=1$.
Hardness results were also proven in FGS18 in relation to workload matrices where only two flow-paths need to be determined. However, the problem in hand is not RRP but a variant of RRP, which we call ${}^{\prime}$, where the objective to be minimized is the maximum workload cost over all demands (rather than the sum).
Theorem 3 ((FGS18, , Theorem 4)).
The problem ${}^{\prime}$($\Delta_{S}=1$) is NP-complete. Moreover, we may assume that instances are such that: each switch has $4$ ports; all entries of the workload matrix are 00 except two which are $1$; and the weight of any dynamic link is $\mu=0$.
The proof of Theorem 3 in FGS18 holds even if the weight of every dynamic link is fixed at $\mu>0$. However, the proof is reliant on there being a static link of weight exponential in $n$ (where $n$ is the number of nodes). In fact, their model is much more permissive than ours and this theorem does not translate into a result about RRP as we have defined it.
The hardness results in FPS19 were primarily concerned with the situation when there is only one switch and every node has the same number of external ports.
Theorem 4 ((FPS19, , Theorems 4.1 and 4.2 and Corollary 4.3)).
The problemsRRP ($\Delta_{S}\geq 2$, $\sigma\geq 0$) is NP-complete. Moreover, we may assume that instances are such that: the static network has no links; there is one switch and every node of the static network is adjacent to the switch via $\Delta_{S}$ external ports; and the weight of any dynamic link is $\mu=1$.
The proof of Theorem 4 in FPS19 does not enable us to restrict the workload matrix in any meaningful way.
We reiterate that Theorems 1-4 all hold for our model within which all dynamic links uniformly have the same weight $\mu$ (as specified in the results).
In contrast to the NP-completeness results above, certain restrictions of RRP were proven in FGS18 ; FPS19 to be solvable in polynomial-time.
Theorem 5 ((FGS18, , Theorem 3)).
The problem RRP can be solved in polynomial-time when the workload matrix has exactly one non-zero entry. Moreover, the problem remains solvable in polynomial-time even if we allow non-uniform weights for dynamic links but so long as the triangle inequality holds for dynamic links.
In Theorem 5, to say that the triangle inequality holds for dynamic links means that whenever $(u,v)$, $(v,w)$ and $(u,w)$ are dynamic links, we have that the weight of $(u,v)$ plus the weight of $(v,w)$ is at least the weight of $(u,w)$.
Theorem 6 ((FPS19, , Theorem 3.1)).
Note that Theorem 6 subsumes (FGS18, , Theorem 1).
An additional complexity-theoretic hardness result was proven in FFS19 ; FFSV20 . However, this result is with respect to a significantly different model.
Theorem 7 ((FFS19, ; FFSV20, , Theorem 1)).
If all links are uni-directional and we allow non-uniform dynamic link weights then the problem RRP cannot be approximated within a ratio of $\Omega(\log n)$ unless ${\bf P}={\bf NP}$.
The proof of Theorem 7 in FFS19 ; FFSV20 is strongly reliant on the more generous model (specifically with dynamic links weights in $[1,100n^{2}+3]$) and it is not obvious that the result holds in our model. Given the hardness result in Theorem 7, the reconfiguration routing problem (as defined in the statement of Theorem 7) was also empirically considered in FFS19 ; FFSV20 with a variety of heuristic algorithms for its solution studied in comparison with an integer linear programming algorithm.
3.2 Motivation for our analysis
The results from the previous section provide various refined versions of the complexity-theoretic hardness of the problem RRP and also some tractability results. However, the refinement scenarios can be improved. Consider the weights of dynamic links. Optical switches provide the primary technological motivation for dynamic links.
Our model insists that all dynamic links have the same weight $\mu$ (with all switch links having weight 00). This is a reasonable assumption as any switch is proprietorial and on such a small scale that surely the internal connections in any switch matching can be taken to have identical weights. Also, we assume that the switch links and the internal link between switch ports that form a dynamic link are of the same communications medium, e.g., optical, and so our consideration of a dynamic link as ‘atomic’ with the overall weight set at $\mu$ is also reasonable. The actual value for $\mu$ will be governed by technology; for example, if dynamic links are optical and static links electrical then the value of $\mu$ will reflect the different transmission costs over these media. An ideal scenario is for our complexity results to hold no matter what the specific value for $\mu$ (in relation to the weights of the static links). So, a constant value for all dynamic link weights is more realistic. Moreover, bearing in mind our technological motivation, the actual value of $\mu$ should be much less than the (normalized) cost of $1$ for a static link (and preferably hardness results should hold for a range of values for $\mu$). Whereas $\mu$ is indeed constant in Theorems 1-4, the actual value of $\mu$ for which three of these results hold is $1$, i.e., the same as the weight of a static link, with the value of $\mu$ equal to 00 for the other result. The proof of Theorem 7 only holds when non-uniform dynamic link weights are allowed (and that result is for a significantly different underlying model). We look at refining hardness results for RRP when the weight of a dynamic link is uniform and more realistic in comparison to the weight of a static link.
The intended domain in which the results of the previous section are to be applied is primarily that of reconfigurable data centre networks. As explained earlier, data centre networks almost always have structure. None of the results from the previous section are specific to a given family of static networks. We look at refining hardness results when the hardness results apply to specific families of static networks that have structure and are closer to the families of interconnection networks that are studied as the communication fabrics of modern-day data centres and supercomputers. This direction for future research was mentioned in the conclusion of FPS19 .

### v2

# Upper Mountjoy Campus, Stockton Road, Durham DH1 3LE, UK

@@ Abstract
A hybrid network is a static (electronic) network that is augmented with optical switches. The Reconfigurable Routing Problem (RRP) in hybrid networks is the problem of finding settings for the optical switches augmenting a static network so as to achieve optimal delivery of some given workload. The problem has previously been studied in various scenarios with both tractability and NP-hardness results obtained. However, the data center and interconnection networks to which the problem is most relevant are almost always such that the static network is highly structured (and often node-symmetric) whereas all previous results assume that the static network can be arbitrary (which makes existing computational hardness results less technologically relevant and also easier to obtain). In this paper, and for the first time, we prove various intractability results for RRP where the underlying static network is highly structured, for example consisting of a hypercube, and also extend some existing tractability results.
Keywords:
algorithms complexity reconfigurable topologies optical circuit switches software-defined networking.
The principal improvement to this paper compared to the previous version is the addition of Subsection 3.1.
@@ 1 Introduction
The rapid growth of cloud computing applications has induced demand for new technologies to optimize the performance of data center networks dealing with ever-larger workloads. The data center topology design problem (that of finding efficient data center topologies) has been studied extensively and resulted in myriad designs (see, e.g., [7]). Advances in hardware, such as optical switches reconfigurable in milli- to micro-seconds, have enabled the development of reconfigurable topologies (see, e.g., [19]). These topologies can adjust in response to demand (demand-aware reconfigurable topologies) or vary configurations over time according to a fixed protocol (demand-oblivious reconfigurable topologies; see, e.g., [4]). So-called hybrid data center networks are a combination of a static topology consisting of, for example, electrical switches, and a demand-aware reconfigurable topology implemented, for example, with optical circuit switches or free space optics (see, e.g., [6, 14, 20, 25]). An intuitive example of a simple reconfigurable topology is illustrated in Fig. 1.
Figure 1: Basic model of an optical wireless data-center network, as described in [6, 20, 25]. Practical timescales for reconfiguration vary from milliseconds [20] to microseconds or nanoseconds [6, 25].
The hybrid network paradigm combines the robustness guarantees of static networks with the ability of demand-aware reconfigurable networks to serve large workloads at very low cost. Consider, for example, the hybrid network shown in Fig. 2, and the configuration shown in Fig. 3. In the (unaugmented) static network, there are two possible paths along which a message from node $b$ to node $d$ may be routed: $b\static f\static h\static e\static d$ or $b\static f\static e\static d$. In the hybrid network as configured in Fig. 3, the path $b\dynamic a\static c\dynamic d$ (among others) is an option [^1] We denote by $u\dynamic v$ the concatenation of a switch link from $u$ to some switch, of the internal switch connection, and of a switch link to $v$ from that switch..
Figure 2: A hybrid network.
Figure 3: An augmented network and its abstracted dynamic links.
@@
Of particular interest to us is the question of how the reconfigurable (optical) portion of the network should be configured for some demand pattern, formalized by Foerster, Ghobadi and Schmid [12] as the Reconfigurable Routing Problem (RRP): in short, given a hybrid network (consisting of a static network and of some switches) and a workload, we wish to choose a configuration (setting of the switches) which results in an optimal delivery of the workload.
@@
Crucially, existing hardness results are only valid when the static network is allowed to be arbitrary, which is almost never the case in practice where interconnection and data center network design is driven by symmetry, high connectivity, recursive decomposition, and so forth. For example: the popular switch-centric data center network Fat-Tree [3] is derived from a folded Clos network; the server-centric data center network DCell [17] is recursively-structured whereby at each level, a graph-theoretic matching of servers is imposed; and the server-centric data center network BCube [16] is recursively-structured with a construction based around a generalized hypercube. (It should be noted that there do exist examples of unstructured data center networks, such as Jellyfish [23] and Xpander [24] which utilize the theory of random graphs.) Many (but not all) NP-complete problems become tractable when the input is restricted to the graphs providing the communications fabric for data center networks and other interconnection networks. For example, Hamiltonian paths are often trivial to find in many interconnection networks; indeed, no finite connected vertex-transitive graph without a Hamiltonian path is known to exist (the Lovász Conjecture contends there is no such graph - see Section 4 of [22]). This motivates our investigation into how the complexity of RRP changes when we restrict to more structured and realistic networks. The question of the complexity of RRP for specific network topologies was specifically identified as an area for future work in [11].
@@
In this paper, we establish for the first time hardness results for RRP that apply to various specific families of highly structured static networks such as, for example, the hypercubes. Our constructions are (perhaps not surprisingly) of a much more involved nature than has hitherto been the case.
2 Problem Setting
The decision problem Reconfigurable Routing Problem considered in this paper is a proper restriction of that presented in prior work [11, 12, 13]. In this section, we provide technical detail to fully formalize our version of the problem, but also additionally provide sufficient framing to briefly review existing results and to identify the areas strengthened by our contribution.
@@ 2 Problem Setting
2.1 Hybrid networks, (re)configurations and (segregated) routing
@@
A hybrid network $G(S)$ can be visualized as in Fig. 2, and consists of a static network $G$ and some switches $S$ augmenting it. A static network $G$ can be abstracted as an undirected graph $G=(V,E)$ so that each static link $(u,v)\in E$ has some fixed weight $w\in\mathbb{Q}_{+}$ (reflecting a transmission cost) and is incident with internal ports of two distinct nodes of $V$. The number of internal ports of some node $v\in V$ is then exactly the degree of $v$ in the abstracted graph $G$. We denote by $S$ a set of switches augmenting the static network $G$ with switch links joining switch ports of some switch to external ports of some of the nodes of $V$. Every switch link has weight 00 (we say more about switch link weights momentarily). Every switch $s\in S$ has at least two switch ports.
In general, the number of external ports of the nodes of a static network $G=(V,E)$ is variable, as is the number of switch ports of the switches of a hybrid network $G(S)$, and it may be the case that there is more than one switch link between a specific node and a specific switch. We assume that the switch links describe a bijection between the external ports and the switch ports; otherwise, there would be some unused ports, which we can safely ignore.
@@
Given a hybrid network $G(S)$ and a switch $s\in S$, a switch matching $N_{s}$ of $s$ is a set of pairs of switch ports of $s$ so that all switch ports involved are distinct. Each switch matching represents an internal setting of the switch and naturally yields a set of pairs of external ports of nodes where all such ports are distinct; we refer to a set of pairs of external ports obtained in this way as a node matching (note that this differs from the standard graph-theoretic notion of a matching). An illustration of a configured hybrid network is shown in Fig. 3: on the left side, switch matchings are represented as sets of arcs, and on the right side the corresponding node matching is shown as a set of dotted lines.
A configuration $N$ is a set of switch matchings, one for each switch. A configuration straightforwardly encodes the corresponding node matchings. We say that $(u,v)$ is a dynamic link in the configuration $N$ (we sometimes write $(u,v)\in N$) if $(u,v)$ appears in any node matching corresponding to $N$.
We allocate a fixed weight $\mu\in\mathbb{Q}_{+}$ to each internal port-to-port connection in a switch $s$. Although a dynamic link is an atomic entity, it can be visualized as consisting of a switch link followed by an internal port-to-port connection in $s$ followed by another switch link. We denote by $G(N)$ the static network $G$ augmented with the dynamic links (each of weight $\mu$) resulting from the configuration $N$ and we call $G(N)$ an augmented network. In the augmented network visualized in Fig. 3, for example: $(a,b)$ is a dynamic link; $(a,c)$ is a static link; and $(e,h)$ is both a static link a dynamic link. Note that it is possible that an augmented network $G(N)$ is a multigraph.
@@
The concepts defined above are driven by reconfigurable hardware technology such as optical switches, wireless (beamforming) and free-space optics, all of which establish port-to-port connections, i.e., switch matchings. The survey paper [14] provides some detail as regards the relationship between the emergent theoretical models and current opto-electronic technology.
2.2 Routing in hybrid networks
Consider again the example shown in Fig. 3. In the configuration shown, a message $M$ from $c$ to node $e$ may be routed:
  1. 1.
    via static links only, along the path $\varphi_{1}:=c\static b\static f\static e$ with weight $3w$, or
  2. 2.
    via dynamic links only, along the path $\varphi_{2}:=c\dynamic d\dynamic h\dynamic e$ with weight $3\mu$, or
  3. 3.
    via a combination of static and dynamic links, along the path $\varphi_{3}:=c\dynamic d\static e$ with weight $\mu+w$.
Depending on the value of $\mu$, any of the paths may minimize the cost to route $M$: if $\mu\geq 2w$ then $\varphi_{1}$ is optimal; if $\mu\leq\frac{w}{2}$ then $\varphi_{2}$ is optimal; and if $\mu\in[\frac{w}{2},2w]$ then $\varphi_{3}$ is an optimal. We may wish to bound the number of alternations allowed between optic and static links in any path a message takes; we capture this hardware requirement via a segregation parameter $\sigma\in\mathbb{N}\cup\{\infty\}$, as introduced in [13], that is the number of alternations between static and dynamic links. In the fully segregated case, $\sigma=0$: messages may be routed either by static links only (as in $\varphi_{1}$) or by dynamic links only (as in $\varphi_{2}$). In the non-segregated case, $\sigma=\infty$ and there is no restriction on the number of alternations, so any path is admitted. Note $\varphi_{3}$ is admitted as a valid path to route $M$ if and only if $\sigma\geq 1$. The dynamic link limit $\delta$, like the segregation parameter $\sigma$, is a restriction on admissible flow-paths. Whereas $\sigma$ describes the maximum number of alternations between static and dynamic links permitted, $\delta$ describes the maximum number of dynamic links any flow-path may use. In particular, when $\delta=1$ every flow-path must contain at most one dynamic link.
Networks are expected to route many messages (of varying sizes) optimally at the same time. Given a hybrid network $G(S)$ we represent the set of all demands we must optimize for as a workload (matrix) $D$ with entries $\{D[u,v]\in\mathbb{Q}_{+}:u,v\in V\}$ providing the intended pairwise node-to-node workloads (each $D[u,u]$ is necessarily 00).
@@ 2.2 Routing in hybrid networks
Given a configuration $N$ and $u,v\in V$ for which $D[u,v]>0$, we route the corresponding workload via a path in $G(N)$ from $u$ to $v$ in $G(N)$ so that this chosen flow-path $\varphi(u,v)$ has workload cost $D[u,v]\times wt_{G(N)}(\varphi(u,v))$, where the weight $wt_{G(N)}(\varphi(u,v))$ is the sum of the weights of the links of the flow-path $\varphi(u,v)$ (if $G(N)$ has both a static link $(x,y)$ and a dynamic link $(x,y)$ then we need to say which we are using in $\varphi(u,v)$). The total workload cost (of $D$ under $N$) is defined as
@@
Our aim will be to find a configuration $N$ in some hybrid network $G(S)$ and flow-paths in $G(N)$ for which the total workload cost of some workload matrix $D$ is minimized. In an unrestricted scenario, we would choose any flow-path $\varphi(u,v)$ to be a flow-path of minimum weight from $u$ to $v$ in $G(N)$, the weight of which we denote by $wt_{G(N)}(u,v)$. When $\sigma\neq\infty$ we must also ensure the flow-path has at most $\sigma$ alternations. We also have the analogous concepts $wt_{G}(\varphi(u,v))$ and $wt_{G}(u,v)$ where we work entirely in the static network $G$. Note that we often describe $D$ by a weighted digraph, which we usually call $D^{\prime}$, so that the node set is $V$ and there is an edge $(u,v)$ of weight $w>0$ if, and only if, $D[u,v]=w$. We also refer to some $D[u,v]>0$ as a demand (from $u$ to $v$).
2.3 The Reconfigurable Routing Problem
We are now in a position to introduce our protagonist:
Reconfigurable Routing Problem ($\sigma$) (RRP($\sigma$))
Input:
$(G,S,\mu,w,D,\kappa)$: $D$ is a workload matrix for the hybrid network $G(S)$ with static (resp. dynamic) links all of weight $w$ (resp. $\mu$).
Question: Does $G(S)$ admit some configuration $N$ such that the total workload cost of $D$ under $N$ (where the number of alternations for any path is bounded by $\sigma$) is at most $\kappa$?
As previously alluded to, this setting is more expressive than we require for most of this paper, and more restrictive than the exact formalism considered in prior work [11, 12, 13]: in those works, $w$ and $\mu$ are sometimes allowed to be functions of their endpoints rather than fixed constants. This provides much more expressivity; notably, their model loses no power when it is restricted to inputs where $G$ is a complete graph and there is only one switch, since it is possible to simulate any other instance by assigning prohibitively large weights to any static edges and any pair of switch ports which should not be usable.
We now turn to the “realistic” networks we mentioned in our introduction. Henceforth unless otherwise specified, static link weights are all equal (and normalized to $1$) and dynamic link weights are always some fixed constant $\mu\in\mathbb{Q}_{+}$. Also, there is a single switch and all nodes are connected to it with identical hardware. This is both practically relevant and intuitively realistic; see e.g. Fig. 1. Then the set of switches $S$ of the hybrid network consists of just one switch, which is fully described by the number of switch links each node in the hybrid network has, which we call $\mathrm{\Delta}_{S}$. This is closely related to the maximum reconfigurable degree $\mathrm{\Delta}_{R}$ from [13], which is an upper bound on the number of external ports per node. The resulting restriction of RRP can be formalized as follows:
$\mathrm{\Delta}_{S}$-switched RRP ($\sigma$)
Input:
$(G,\mu,D,\kappa)$: $D$ is a workload matrix for the hybrid network $G(S)$ with static (resp. dynamic) links all have weight $1$ (resp. $\mu$) (where $S$ consists of a single switch that every node in $G$ is connected to exactly $\mathrm{\Delta}_{S}$ times).
Question: Does $G(S)$ admit some configuration $N$ such that the total workload cost of $D$ under $N$ (where the number of alternations for any path is bounded by $\sigma$) is at most $\kappa$?